OpenStudy (anonymous):

You would like to have $1,000,000 in 33 years by making regular deposits at the end of each month in an annuity that pays 7% compounded monthly. Determine the deposit at the end of each month. Do not round until the final answer. Then round to the nearest dollar as needed.

OpenStudy (anonymous):

FV of annuity = payment*[(1+r)^n - 1]/r annual r = 7% monthly r = (1+7%)^(1/12) - 1 = 0.5654% n = 33 years = 33*12 = 396 Monthly payment required = 1000000*0.5654%/[(1+0.5654%)^396 - 1] = 679.1749946 = $679 (rounded)

OpenStudy (anonymous):

Future value= \[ \frac{R \left((i+1)^n-1\right)}{i} \] Where R is the monthly payment \[ \frac{R \left((\frac {.07}{12}+1)^{396}-1\right)}{\frac {.07}{12}}=10^6 \] Solve for R and you get R=647.64

OpenStudy (anonymous):

@smicatrotto r=.07/12

OpenStudy (anonymous):

Oh okay, thank you!!